# Learning Actionable Counterfactual Explanations in Large State Spaces

### v2

@@ Abstract
Recourse generators provide actionable insights, often through feature-based counterfactual explanations (CFEs), to help negatively classified individuals understand how to adjust their input features to achieve a positive classification. These feature-based CFEs, which we refer to as low-level CFEs, are overly specific (e.g., coding experience: $4\to 5+$ years) and often recommended in feature space that doesn’t straightforwardly align with real-world actions. To bridge this gap, we introduce three novel recourse types grounded in real-world actions: high-level continuous (hl-continuous), high-level discrete (hl-discrete), and high-level ID (hl-id) CFEs.
We formulate single-agent CFE generation methods, where we model the hl-discrete CFE as a solution to a weighted set cover problem and the hl-continuous CFE as a solution to an integer linear program. Since these methods require costly optimization per agent, we propose data-driven CFE generation approaches that, given instances of agents and their optimal CFEs, learn a CFE generator that quickly provides optimal CFEs for new agents. This approach, also viewed as one of learning an optimal policy in a family of large but deterministic MDPs, considers several problem formulations, including formulations in which the actions and their effects are unknown, and therefore addresses informational and computational challenges.
Through extensive empirical evaluation using publicly available healthcare datasets (BRFSS, Foods, and NHANES), we compare the proposed forms of recourse to low-level CFEs and assess the effectiveness of our data-driven approaches. Empirical results show that the proposed data-driven CFE generators are accurate and resource-efficient, and the proposed forms of recourse have various advantages over the low-level CFEs.
@@ 1 Introduction
Machine learning models are increasingly being used to guide high-stakes decision-making processes. Given the potential impact on individuals’ livelihoods, society demands transparency and the right to an explanation, as outlined in Articles 13–15 of the EuropeanParliament2016a General Data Protection Regulation. A critical aspect of this transparency is understanding how individuals (agents) can modify their input features to achieve a desired outcome, such as a positive label in a binary classification setting. Recourse generators that provide actionable insights offer one such solution (Wachter2017; Dandl2020; Mothilal20; Ustun19; Karimi21; Shalmali19; Karimi22).
A popular form of recourse, actionable recourse Ustun19, generates feature-based counterfactual explanations (CFEs), specifying precise adjustments to features to ensure that the new features collectively result in a positive classification. For comparison with our work, we refer to these as low-level CFEs. While helpful, low-level CFEs are overly specific and might be challenging to translate into real-world-like actions. To address this limitation, we introduce three novel forms of recourse that align with real-world actions: high-level continuous (hl-continuous), high-level discrete (hl-discrete), and high-level ID (hl-id) CFEs.
@@
Figure 1 presents an illustrative example of hypothetical recourse across three binary classification tasks, highlighting the distinctions between low-level and high-level CFEs. Specifically, the hl-continuous CFEs involve general and predefined real-world-like actions, simultaneously modifying multiple decision-making features through numerical adjustments. Similarly, hl-discrete CFEs also stem from real-world actions and might modify several features simultaneously. However, unlike hl-continuous CFEs, which adjust feature values, hl-discrete CFEs modify feature eligibility. The hl-discrete CFEs use binary vector actions to ensure all features meet a predefined threshold, reducing decisions to unit simple yes/no questions. It is particularly efficient in scenarios where feature satisfiability depends on a set threshold, such as level-one decision-making in wellness checks (see Figure 1). Lastly, the hl-id CFE encapsulates either hl-continuous or hl-discrete CFEs into a single overarching action, encompassing all necessary changes without detailing specific feature modifications. This abstraction often conveys significant implicit information.
@@
Finally, we conduct extensive experiments on $11+$ agent–CFE datasets derived from real-world healthcare datasets: BRFSS, Foods, and NHANES. We chose these datasets due to the availability of a wealth of publicly accessible data, which allows us to effectively demonstrate the benefits of incorporating real-world-like actions in CFE generation and sufficiently explore the impact of a larger action space (or action grid according to Ustun19), enabled by the high number of actionable features with broad value ranges. Alongside these real-world datasets, we also include $11+$ fully-synthetic agent–CFE datasets for comparison. We extensively compare the low-level CFEs with the hl-continuous and hl-discrete CFEs, and provide an in-depth analysis of the performance of the data-driven hl-continuous, hl-discrete, and hl-id CFE generators under various settings (Sections LABEL:sec:exp_setup and LABEL:sec:results). Our code can be accessed here: link.
@@ 2 Background
We consider a binary classification setting, where an agent with state $\mathbf{x}$ receives either a positive (desirable) or negative (undesirable) classification under a model $f(\mathbf{x})$. Although we focus on this setting, our proposed data-driven CFE generation framework generalizes to other scenarios. Given an agent state $\mathbf{x}$ with an undesirable model outcome, the objective of the CFE generator is to provide the agent with information that they can act on to achieve a desirable classification under the model. Actionable recourse (Ustun19), generates low-level CFEs where each action in the CFE precisely specifies how much the agent should add or subtract from a specific feature to ensure that, collectively, the new features (state) result in the agent receiving a desirable model outcome.
@@
Shortcomings of low-level CFEs and their generators. We note two limitations of low-level CFEs in comparison to our proposed forms of recourse (hl-continuous, hl-discrete, and hl-id CFE), and two, as we compare the low-level CFE generator to the proposed data-driven CFE generation approach (Verma2020CounterfactualEF; Karimi22; hiddenSolon).
First, low-level CFEs are feature-based and highly specific (e.g., in Figure 1, app1, carbohydrate (gm): $109.45\to 43.376000000000005$), which may overwhelm agents and introduce additional costs to translate the CFE into implementable steps. In contrast, our proposed CFEs are better aligned with real-world scenarios, offering what you see is what you get actionable insights (see Figure 1). Furthermore, with low-level CFEs, details about the actions the agent implements (the number of them needed, which features they would simultaneously modify, and costs to incur) are often unknown beforehand, potentially leading to a misleading price of recourse and related metrics such as sparsity (few modified features) and proximity (closeness of final state to initial state) (hiddenSolon).
Second, although a CFE (e.g., complete the Google AI residency program in Figure 1) could have been optimal for several agents with different but close profiles (e.g., one has coding experience: $4$ and another $3$), the low-level CFE being too specific, would give the agents different CFEs (coding experience: $4\to 5+$ years and coding experience: $3\to 5+$ years). In contrast, our proposed recourse methods are both agent-specific (tailored to an agent’s initial state) and generalizable (providing similar recommendations to agents with comparable profiles).
Lastly, low-level CFE generators operate on a single-agent basis (Verma2020CounterfactualEF; Karimi22), making them relatively expensive. In contrast, once trained, our proposed data-driven approach can generate CFEs for multiple new agents without the need for generator re-optimization. Moreover, low-level CFE generation becomes very limited, sometimes impractical, in scenarios where access to critical information, such as sufficient training data, classifier parameters, or a comprehensive list of actions and their costs, is restricted or inaccessible. Our proposed data-driven CFE generation would work better in this case, provided there is access to historical mappings of agents and their optimal CFEs, e.g., healthcare interventional data or high school counselors’ past successful recommendations for college applicants.
@@ 3 The Proposed Single-agent CFE Generators
This section outlines the single-agent CFE generators for the proposed hl-continuous and hl-discrete CFEs. Each generator relies on predefined, real-world-like actions to solve optimization problems for CFE generation: the weighted set cover problem for hl-discrete CFEs and integer linear programming for hl-continuous CFEs.
@@ 3.1 The Single-agent hl-continuous CFE Generation
Below, we formally define hl-continuous actions and the single-agent hl-continuous CFE generation process.
@@ 3.2 The Single-agent hl-discrete CFE Generation
Below, we formally define hl-discrete actions and the single-agent hl-discrete CFE generation process.

### v3

@@ Abstract
Recourse generators provide actionable insights, often through feature-based counterfactual explanations (CFEs), to help negatively classified individuals understand how to adjust their input features to achieve a positive classification. These feature-based CFEs, which we refer to as low-level CFEs, are overly specific (e.g., coding experience: $4\to 5+$ years) and often recommended in a feature space that doesn’t straightforwardly align with real-world actions. To bridge this gap, we introduce three novel recourse types grounded in real-world actions: high-level continuous (hl-continuous), high-level discrete (hl-discrete), and high-level ID (hl-id) CFEs.
We formulate single-agent CFE generation methods for hl-discrete and hl-continuous CFEs. For the hl-discrete CFE, we cast the task as a weighted set cover problem that selects the least cost set of hl-discrete actions that satisfy the eligibility of features, and model the hl-continuous CFE as a solution to an integer linear program that identifies the least cost set of hl-continuous actions capable of favorably altering the prediction of a linear classifier. Since these methods require costly optimization per agent, we propose data-driven CFE generation approaches that, given instances of agents and their optimal CFEs, learn a CFE generator that quickly provides optimal CFEs for new agents. This approach, also viewed as one of learning an optimal policy in a family of large but deterministic MDPs, considers several problem formulations, including formulations in which the actions and their effects are unknown, and therefore addresses informational and computational challenges.
We conduct extensive empirical evaluations using publicly available healthcare datasets (BRFSS, Foods, and NHANES) and fully-synthetic data. For negatively classified agents identified by linear and threshold-based binary classifiers, we compare the proposed forms of recourse to low-level CFEs, which suggest how the agent can transition from state $\mathbf{x}$ to a new state $\mathbf{x}^{\prime}$ where the model prediction is desirable. We also extensively evaluate the effectiveness of our neural network-based, data-driven CFE generation approaches. Empirical results show that the proposed data-driven CFE generators are accurate and resource-efficient, and the proposed forms of recourse offer various advantages over the low-level CFEs.
@@ 1 Introduction
Machine learning models are increasingly being used to guide high-stakes decision-making processes. Given the potential impact on individuals’ livelihoods, society demands transparency and the right to an explanation, as outlined in Articles 13–15 of the EuropeanParliament2016a General Data Protection Regulation and Article 13 of the EuropeanParliament2025 AI Act. A critical aspect of this transparency is understanding how individuals (agents) can modify their input features to achieve a desired outcome, such as a positive label in a binary classification setting. Recourse or counterfactual explanation (CFE) generators that provide actionable insights offer one such solution (Wachter2017; Ustun19; Shalmali19; Dandl2020; Mothilal20; Karimi21; Karimi22).111 Following prior work  (pawelczyk2022exploring; Rasouli24; Jiang2024; Verma2020CounterfactualEF), we use the terms recourse and CFE interchangeably. For a detailed discussion and a nuanced differentiation of these terms, we refer the reader to Section 2 of Karimi22 and Section 3.3 of Verma2020CounterfactualEF.
A popular form of recourse generation, actionable recourse (Ustun19), generates feature-based CFEs, specifying precise adjustments to features (state) to ensure that the new features collectively result in a positive classification. For comparison with our work, we refer to these as low-level CFEs. While helpful, low-level CFEs are overly specific and might be challenging to translate into real-world-like actions. To address this limitation, we introduce three novel forms of recourse that align with real-world actions: high-level continuous (hl-continuous), high-level discrete (hl-discrete), and high-level ID (hl-id) CFEs.
@@
Figure 1 presents an illustrative example of hypothetical recourse across three binary classification tasks, highlighting the distinctions between low- and high-level CFEs. Specifically, the hl-continuous CFEs involve general and predefined real-world-like actions that modify multiple features simultaneously through numerical adjustments. Similarly, hl-discrete CFEs also stem from real-world actions, each of which might affect several features, with the effect on each feature assumed to be known. However, unlike hl-continuous CFEs, which adjust feature values, hl-discrete CFEs modify feature eligibility. The hl-discrete CFEs use binary vector actions to ensure all features meet a predefined threshold, reducing decisions to unit simple yes/no questions. It is particularly efficient in scenarios where feature satisfiability depends on a set threshold, such as level-one decision-making in wellness checks (see Figure 1 and Appendix LABEL:sec:app_hld-hlc-diff). Lastly, the hl-id CFE encapsulates either hl-continuous or hl-discrete CFEs into a single overarching action, encompassing all necessary changes without detailing specific feature modifications. This level of abstraction relies on domain knowledge and often conveys significant implicit information.
@@
Finally, we conduct extensive experiments on $30$ agent–CFE datasets derived from real-world healthcare datasets: BRFSS, Foods, and NHANES. We chose these datasets due to the availability of a wealth of publicly accessible data, which allows us to effectively demonstrate the benefits of incorporating real-world-like actions in CFE generation and sufficiently explore the impact of a larger action space (or action grid according to Ustun19), enabled by the high number of actionable features with broad value ranges. Alongside these real-world datasets, we also include $34$ fully-synthetic agent–CFE datasets for comparison. We extensively compare the low-level CFEs with the hl-continuous and hl-discrete CFEs, and provide an in-depth analysis of the performance of the data-driven hl-continuous, hl-discrete, and hl-id CFE generators under various settings (Sections LABEL:sec:exp_setup and LABEL:sec:results). Our code can be accessed here.
@@ 2 Background
We consider a binary classification setting, where an agent in the state $\mathbf{x}\in\mathcal{X}$ receives either a positive (desirable) or negative (undesirable) outcome under a model $f(\mathbf{x})$. The state space $\mathcal{X}$ consists of all valid agent states, each represented by a feature vector capturing attributes such as age and calcium(mg). The model operates over this space, and the CFE generator searches within it to identify alternative states with favorable model outcomes. Although we focus on binary classification, our data-driven CFE framework generalizes to other settings. Given an undesirable outcome, the CFE generator suggests actionable changes to move the agent at state $\mathbf{x}$ to a new state $\mathbf{x}^{\prime}\in\mathcal{X}$ where the model prediction is desirable. Actionable recourse (Ustun19) provides low-level CFEs that specify feature-level changes needed to reach such a state.
@@
Shortcomings of low-level CFEs and their generators. We note two limitations of low-level CFEs in comparison to our proposed forms of recourse (hl-continuous, hl-discrete, and hl-id CFE), and two, as we compare the low-level CFE generator to the proposed data-driven CFE generation approach (hiddenSolon; Karimi22; Verma2020CounterfactualEF).
First, low-level CFEs are feature-based and highly specific (e.g., in Figure 1, app1, calcium (mg): $309\to 113$), which may overwhelm agents and introduce additional costs to translate the CFE into implementable steps. In contrast, our proposed CFEs are better aligned with real-world scenarios, offering what you see is what you get actionable insights (see Figure 1). Furthermore, with low-level CFEs, details about the actions the agent implements (the number of them needed, which features they would simultaneously modify, and costs to incur) are often unknown beforehand, potentially leading to a misleading price of recourse and related metrics such as sparsity (few modified features) and proximity (closeness of final state to initial state) (hiddenSolon).
Second, although a CFE (e.g., complete the Google AI residency program in Figure 1) could have been optimal for several agents with different but close profiles (e.g., one has coding experience: $4$ and another $3$), the low-level CFE being too specific, would give the agents different CFEs (coding experience: $4\to 5+$ years and coding experience: $3\to 5+$ years). In contrast, our proposed recourse generation approaches are both agent-specific (tailored to an agent’s initial state) and generalizable (providing similar recommendations to agents with comparable profiles).
Lastly, while most low-level CFE generators operate on a single-agent basis (Karimi22; Verma2020CounterfactualEF), recent work by Pedapati2020; Rawal2020; pmlr-v151-kanamori22a; Ley2023 and Carrizosa24 propose global low-level CFE generators capable of producing CFEs for multiple agents. However, these methods generate feature-based CFEs and require, at a minimum, query access to the classifier. In contrast, we propose data-driven CFE generation approaches that, given historical mappings between agents and their optimal CFEs, such as healthcare intervention records or high school counselors’ past successful college recommendations, can learn to generate CFEs with real-world-like actions for multiple new agents without requiring re-optimization. Additionally, the proposed data-driven approaches work well in settings where access to critical information, such as sufficient classification training data, classifier, or a comprehensive list of actions and their costs, is restricted or inaccessible.
@@ 3 The Proposed Single-agent CFE Generators
This section outlines the single-agent CFE generators for the proposed hl-continuous and hl-discrete CFEs. Each generator relies on predefined, real-world-like actions to solve optimization problems for CFE generation: the weighted set cover problem for hl-discrete CFEs and ILP for hl-continuous CFEs.
@@ 3.1 The Single-agent hl-continuous CFE Generation
Below, we formally define hl-continuous actions and the single-agent hl-continuous CFE generation process for outcomes predicted by linear classification models.
@@ 3.2 The Single-agent hl-discrete CFE Generation
Below, we formally define hl-discrete actions and the single-agent hl-discrete CFE generation process based on a threshold classifier that determines feature eligibility.